OpenStudy (anonymous):

Solve the following formula for t. J=W(1+rt)

OpenStudy (anonymous):

You isolate the t using the same method as you would if the non-t variables were constants.

OpenStudy (anonymous):

do you have to distribute first?

OpenStudy (anonymous):

You could but it is better to not.

OpenStudy (anonymous):

i'm just confused on how to even get started, i did other problems but the parenthesis on this one is confusing me.

OpenStudy (anonymous):

Kuromeru can you answer my first question i posted please?

OpenStudy (anonymous):

You can look at 1+rt as a variable itself, so let's temporarily call that z. How would you isolate t (which is expressed in our variable z)? So that J=Wz, where z=1+rt. After you isolate z substitute z for what we substituted for z: 1+rt, then continue to solve as usual. Parentheses often denote multiplication so that Wz is the same as saying W(z).

OpenStudy (anonymous):

i'm still completely lost, but thanks for the help.

OpenStudy (anonymous):

JW-1/r=t <--is that the answer?

OpenStudy (anonymous):

No, the answer is \[t=\frac{ J-W }{ Wr }\] You can start by dividing the W from both sides to leave 1+rt = J/W OR you could distribute the W through and solve from there. Distributing from the start in this case would be easier.